OpenStudy (anonymous):

4. What kind of triangle is made by connecting the points A(0, –6), B(3, –6), and C(3, –2)? (1 point) equilateral right isosceles right and isosceles 5. What type of quadrilateral is formed by connecting the points (0, 9), (3, 6), (0, 1), and (–3, 6)? (1 point) rhombus trapezoid kite quadrilateral

OpenStudy (dumbcow):

use distance formula AB = sqrt(9 +0) = 3 AC = sqrt(9+16) = 5 BC = sqrt(0+16) = 4 3^2 + 4^2 = 5^2 9+16 = 25 True Therefore it is a Right Triangle

OpenStudy (anonymous):

what about the second one

OpenStudy (anonymous):

i go to connections too im doin the same thing

OpenStudy (dumbcow):

 it looks like a kite to me :)